A small delivery truck was purchased on January 1 at a cost of $25,000. It has an estimated useful life of four years and an estimated salvage value of $5,000. Prepare a depreciation schedule showing the depreciation expense, accumulated depreciation, and book value for each year under the sum of years digit method. Accum. depr. end of Yr. 2:$14,000
Answer #1
CALCULATION OF THE DEPRECIATION AS PER SUM OF DIGITS METHOD
Purchase Cost of Machine $                      25,000
Less: Salvage Value $                        5,000
Net Value for Depreciation $                      20,000
SUM OF DIGITS = 1+2+3+4 = 10 Gigits
Depreciation for the year 1 = $ 20,000 * 4 / 10 = $                        8,000
Depreciation for the year 2 = $ 20,000 * 3 / 10 = $                        6,000
Depreciation for the year 3 = $ 20,000 * 2 / 10 = $                        4,000
Depreciation for the year 3 = $ 20,000 * 1 / 10 = $                        2,000
DEPRECIATION SCHEDULE
Year Depreciation Amount Accumulated Depreciation Purchase Price Book Value
1 Depreciation for the year 1 $                        8,000 $                8,000 $              25,000 $                17,000
2 Depreciation for the year 2 $                        6,000 $             14,000 $              25,000 $                11,000
3 Depreciation for the year 3 $                        4,000 $             18,000 $              25,000 $                   7,000
4 Depreciation for the year 4 $                        2,000 $             20,000 $              25,000 $                   5,000